Pixel:

Quick Fix Inc. repairs bikes. Their revenue, in dollars, can be modeled by the equation y = 400 + 220x, where x is the number of hours spent repairing bikes. Their overhead cost, in dollars, can be modeled by the equation y=20x^2+160 , where x is the number of hours spent repairing bikes. After how many hours does the company break even? Note: The phrase break even refers to the value where the two functions are equivalent

Pixel:

Im thinking 30

Pixel:

@dude

Pixel:

3,760 overhead cost if x=30

Pixel:

7k revenue x=30

dude:

Just set them equal to each other and solve for x \(400+220x=20x^2+160\)

Pixel:

x=0?

Pixel:

set them to x =0

dude:

No no, just isolate x

Pixel:

OH

Pixel:

i see what you did

Pixel:

400=20x^2-60

Pixel:

400=20x-60 460=20x

Pixel:

23=x

dude:

Step 1: \(400+220x=20x^2+160\) Step 2: \(400+220x\color{red}{-400}=20x^2+160\color{red}{-400}\) Step 3: \(220x\color{red}{-220x}=20x^2-240\color{red}{-220x}\) Solve: \(0=20x^2-220x-240\) From here we can use the quadratic formula

Pixel:

uhhhhh

Pixel:

I kinda maybe sorta didnt learn the formula ._.

KyroKazimake:

You don't have to use it

KyroKazimake:

Every number in the equation is divisible by 20, so just simplify and then factor like you normally would

Pixel:

20+11x=x^2+8

KyroKazimake:

Ok so now you need to put everything on one side. Basically make it look like a normal quadratic equation

Pixel:

y=x^2+11x+28

Pixel:

?

KyroKazimake:

One sec let me write it out

KyroKazimake:



Pixel:

cant you just type it? no offense. But i could read it easier

KyroKazimake:

I can try lol 20+11x=x^2+8 -20 -11x x^2-11x-12=0

Pixel:

and i can just solve it?

Pixel:

Oh it is 12 ;o

Pixel:

Thanks :D

KyroKazimake:

Np :)